Ncert Solutions Maths class 12th

Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th

  • 2.5k

    Questions

  • 0

    Discussions

  • 1

    Active Users

  • 65

    Followers

Have a question about Ncert Solutions Maths class 12th? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

44. Let, E1: event that outcome on the die is 5 or 6

E2: event that outcome on the die is 1,2,3 or 4

A: event of getting exactly one head

P(E1)=26=13

P(E2)=46=23

P(A/E1)= probability of getting exactly one head by tossing the coin three times if she gets 5 or 6 =38

P(A/E2)= probabilit

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

43. Let, E1: event that first group will win

E2: event that second group will win

A: event that new product will produce

P(E1)=0.6=610=35

P(E2)=0.4=410=25

P(A/E1)=P (introducing new product by group A ) =0.7=710

P(A/E2)=P ( introducing new product by group B ) =0.3=310

Therefore, by Baye’s theorem,

P(E2/A)= probability that n

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

42. Let, E1: event items produced by A

E2: event items produced by B

X: event that produced item was found to be defective

P(E1)=60%=35

P(E2)=40%=25

P(X/E1)=P (items produced by machine A which is defective) =2%=2100

P(X/E2)=P ( items produced by machine B which is defective) =1%=1100

Therefore, by Baye’s

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

41. Let, E1: event that the driver is scooter driver

E2: event that the driver is car driver

E3: event that the driver is truck driver

A: event the person meets with accident

Total number of drivers =2000+4000+6000=12000

P(E1)=P (driver is a scooter driver) =200012000=16

P(E2)=P (driver is a car driver) =400012000=13

P(E3)=P (dri

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

40. Let, E1: event of choosing Q headed coin

E2: event of choosing biased coin

E3: event of choosing unbiased coin

P(E1)=P(E2)=P(E3)=13

Let A be the event that shows head,

P(A/E1)=P (coin shows heads, given that it is a headed coin) =1

P(A/E2)=P (coin showing up head, given that it is biased coin

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

39. Let, E1: event person has disease

E2: event person has no disease

A: be event that blood test is positive

As E1 and E2 are events which are complementary to each other,

Then, P(E1)+P(E2)=1

P(E2)=1−P(E1)

P(E1)=0.1%=0.1100=0.001

∴ P(E2)=1−0.001

=0.999

P(A/E1)=P (result is positive given that person has disease) =99%=0.99

P(A/E2)=P (result is positiv

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

38. Let, E1 :event that knows the answer

E2 :event that he guesses the answer

P(E1)=34 and P(E2)=14

Let A be the event that the answer is correct

Also, P(A/E1)=P (correct answer given that he knows) =1

P(A/E2)=P (correct answer given that he guess) =14

Now, probability that he knows the an

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

37. Let, E1 :event that the student in hostel

E2 : event that the student is day scholar

A : be the event of the chosen student get grade A

P(E1)=60%=60100=0.6

P(E2)=40%=40100=0.4

P(A/E1)=P (student getting an A grade is hostler) =30% =30100=0.3

P(A/E2)=P ( student getting an A grade is a day scholar) =20%=20100=0.2

Then, by Ba

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

36. Let E1 be the event of selecting first bag

E2 be the event of selecting second bag

∴P(E1)=P(E2)=12

Let A be the event of getting a red ball

P(A/E1)=P (drawing a red ball from first bag)

=48=12

P(A/E2)=P (drawing a red ball from second bag)

=28=14

Then, probability of drawing a ball from the first bag which is red, is

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

35. Number of balls contain in win =5 red

=5 black

Let the red ball be drawn in the first attempt

∴ P (drawing a red ball) =510=12

By question, if added two red balls to the win, then 7 red balls and 5 black balls contain.

P (drawing a red ball) =712

Let a black ball be

...Read more

Thumbs Up IconUpvote Thumbs Down Icon